OpenStudy (anonymous):

c. Write a linear equation for the line. (4,0) (0,5)

Parth (parthkohli):

First determine the slope. \(\Large \color{Black}{\Rightarrow slope = {\Delta y \over \Delta x } = {y_{2} - y_{1} \over x_{2} - x_{1}} }\) Then after you have done that, then determine the y-intercept by substituting and then setting x to 0.

OpenStudy (anonymous):

wat

Parth (parthkohli):

Oops my browser crashed

Parth (parthkohli):

Slope would be 5/-4 So far we know that the equation is y = 5x/-4 + b Let's substitute. 0 = 5(4)/-4 + b 0 = -5 + b b = 5 So the equation is: \(\Huge \color{Black}{\Rightarrow y = {5 \over -4}x + 5 }\)

OpenStudy (anonymous):

and then I solve and get the answer?

Parth (parthkohli):

What? No. It asked for the equation.

OpenStudy (anonymous):

O my bad :)

OpenStudy (turingtest):

What do you need help with @RioAnne ?

OpenStudy (anonymous):

Nothing I got it thanks anyway

OpenStudy (turingtest):

okydoke!